# Tree-Wasserstein Distance for High Dimensional Data with a Latent Feature Hierarchy

### v1

@@ D.1 Implementation Details
Zeisel 3005 cells 7 4000 genes
CBMC 8617 cells 13 500 genes
Datasets. In real-world word-document datasets, we test four word-document benchmarks in Kusner et al. [2015]: (i) the BBCSPORT consisting of 13243 bags of words (BOW) and 517 articles categorized into five sports types, (ii) the TWITTER comprising 6344 BOW and 2176 tweets in three types of sentiment, (iii) the CLASSIC, including 24277 BOW and 4965 academic papers from four publishers, and (iv) the AMAZON containing 42063 BOW and 5600 reviews of four products. The types of documents are used as labels in classification tasks. In scRNA-seq data analysis, we test two scRNA-seq datasets in Dumitrascu et al. [2021]. The first dataset, the Zeisel dataset, focuses on the mouse cortex and hippocampus [Zeisel et al., 2015]. It comprises 4000 gene markers and 3005 single cells, categorized into seven distinct cell types. The second dataset, the CBMC dataset, is derived from a cord blood mononuclear cell study [Stoeckius et al., 2017]. It comprises 500 gene markers and 8617 single cells, which belong to 13 different cell types. The cell types are used as labels in cell classification tasks. We report the statistics of these benchmarks in Tab. LABEL:tab:dataset_statistics.
@@ E.2 Runtime Analysis
In Fig. 4, we present the runtime performance of our TWD in comparison to other TWDs and Wasserstein distances when applied to scRNA-seq and word-document datasets. Although our TWD is not the fastest method, it consistently ranks as the fourth or fifth quickest. It outperforms deep-based methods like HHC-TW and gHHC-TW, as well as pre-trained OT baselines such as GMD for single-cell RNA sequencing datasets and WMD for word-document datasets. The fastest methods identified are TSWD-1, TR-TW, and MST-TW. The work in [Le et al., 2019] reports that the runtime of TSWD is influenced by the number of tree slices used in sampling. Similarly, slice variants SWCTWD and SWQTWD are also affected by the number of slices. Despite being slightly slower compared to TSWD-1, TR-TW, and MST-TW, our TWD demonstrates significant advantages in terms of classification accuracy, as shown in Tab. LABEL:tab:classification. We also report the runtime of the OT distance using the embedding distance between features as a ground distance, denoted as OT-ED. We note that this computational expense is $O(mn^{3}+m^{3}\log m)$. In contrast, our TWD offers a substantially reduced runtime of $O(m^{1.2})$, thereby demonstrating a significant advantage when handling large datasets. Note that the scalability of the proposed TWD is enhanced by utilizing diffusion landmarks [Shen and Wu, 2022], as detailed in App. D.3. This enhancement allows our TWD to achieve computational performance comparable to TSWD-1, TR-TW, and MST-TW, particularly in datasets with a large number of features, such as the word-document datasets shown in Fig. 4.
@@ E.3.4 Exploring Other Hyperbolic Embedding and the Impact of Curvature
We note that other TW baselines could use other two-dimensional Poincaré representations (e.g., Poincaré embedding (PE) [Nickel and Kiela, 2017] or hyperbolic VAE [Nagano et al., 2019]) and decode the hyperbolic embedding into a tree by HHC [Chami et al., 2020]. We denoted these baselines by PE-HHC-TW and HVAE-HHC-TW, respectively, and present their results in Tab. E.3.5. We see in the table that HHC [Chami et al., 2020] applied to a PE [Nickel and Kiela, 2017] or hyperbolic VAE [Nagano et al., 2019] is less effective than our TWD by a large margin. This discrepancy is attributed to the sensitivity of the recovery of the hidden feature hierarchy using the competing methods, a finding that is consistent with previous works [Sala et al., 2018, Sonthalia and Gilbert, 2020, Lin et al., 2023].
@@ E.3.5 Robustness Analysis Under Gaussian Noise
PE-HHC-TW 87.2$\pm$1.2 69.7$\pm$0.9 93.2$\pm$1.1 86.2$\pm$0.9 81.6$\pm$0.9 74.9$\pm$1.0
HVAE-HHC-TW 88.4$\pm$2.5 64.3$\pm$1.7 90.8$\pm$1.5 87.3$\pm$1.7 80.2$\pm$0.9 75.1$\pm$1.0
@@ F.2 Putting the Features on the Leaves of the Tree
Note that the hyperbolic embedding [Lin et al., 2023], which is consistent with other common hyperbolic embedding methods [Chamberlain et al., 2017, Nickel and Kiela, 2017, 2018, Sala et al., 2018], does not restrict all the features to the leaves. Assigning the features to the leaf nodes is only carried out in the tree decoding (Alg. 1), which is in line with other tree decoding algorithms [Sarkar, 2011, Chami et al., 2020, Indyk and Thaper, 2003]. Note that our tree decoding algorithm generates a weighted binary tree. We assert that this does not pose a restriction because any tree can be effectively transformed into a binary tree [Bowditch, 2007, Cormen et al., 2022]. For instance, consider the flexible tree in Fig. 10 (a): [sport[soccer,football,basketball]] (where “sports” is a parent node with three children: “basketball”, “soccer”, and “football”). This tree can be accurately modeled as a (weighted) binary tree as in Fig. 10 (b): [$c$[$b$[$a$[soccer, football], basketball]], sport] (where “sports”, “basketball”, “soccer”, and “football” are all leaves, $a,b,c$ are internal nodes, and the edges encode the same relationships as the flexible tree above).
@@ F.3 Differentiability of Our TWD
    Our TWD is differentiable (details in App. F.3), which is an important property that is lacking in existing TW baselines.
In addition, we present favorable empirical results in Sec. 5 compared to several TW and Wasserstein distances.

### v2

@@ D.1 Implementation Details
Zeisel 3005 cells 47 4000 genes
CBMC 8617 cells 56 500 genes
Datasets. In real-world word-document datasets, we test four word-document benchmarks in Kusner et al. [2015]: (i) the BBCSPORT consisting of 13243 bags of words (BOW) and 517 articles categorized into five sports types, (ii) the TWITTER comprising 6344 BOW and 2176 tweets in three types of sentiment, (iii) the CLASSIC, including 24277 BOW and 4965 academic papers from four publishers, and (iv) the AMAZON containing 42063 BOW and 5600 reviews of four products. The Word2Vec embedding [Mikolov et al., 2013] is used as the word embedding vectors that are trained on the Google News [^1] https://code.google.com/archive/p/word2vec/ dataset. The types of documents are used as labels in classification tasks. In scRNA-seq data analysis, we test two scRNA-seq datasets in Dumitrascu et al. [2021], where the datasets are available at the link [^2] https://github.com/solevillar/scGeneFit-python/tree/62f88ef0765b3883f592031ca593ec79679a52b4/scGeneFit/data_files. The first dataset, the Zeisel dataset, focuses on the mouse cortex and hippocampus [Zeisel et al., 2015]. It comprises 4000 gene markers and 3005 single cells. The second dataset, the CBMC dataset, is derived from a cord blood mononuclear cell study [Stoeckius et al., 2017]. It comprises 500 gene markers and 8617 single cells. We apply the divisive biclustering method from Zeisel et al. [2015] to obtain 47 subclasses for the Zeisel dataset and 56 subclasses for the CBMC dataset. The cell subclasses are used as labels in cell classification tasks. The Gene2Vec embeddings [Du et al., 2019] are used as the gene embedding vectors following the study of Gene Mover’s Distance [Bellazzi et al., 2021].
We report the statistics of these benchmarks in Tab. LABEL:tab:dataset_statistics.
Baselines. We compare the proposed TWD with existing TWDs, including (i) Quadtree [Indyk and Thaper, 2003] that constructs a random tree by a recursive division of hypercubes, (ii) Flowtree [Backurs et al., 2020] that computes the optimal flow of Quadtree, (iii) tree-sliced Wasserstein distance (TSWD) [Le et al., 2019] that computes an average of TWDs by random sampling trees, where the numbers of sampling are set to 1, 5, and 10, (iv) UltraTree [Chen et al., 2024] that constructs an ultrametric tree [Johnson, 1967] by minimizing OT regression cost, (v) weighted cluster TWD (WCTWD) [Yamada et al., 2022], (vi) weighted Quadtree TWD (WQTWD) [Yamada et al., 2022], and their sliced variants (vii) SWCTWD and (viii) SWQTWD [Yamada et al., 2022]. We also compare to TWDs where the trees are constructed by (i) MST [Prim, 1957], (ii) Tree Representation (TR) [Sonthalia and Gilbert, 2020] through a divide-and-conquer approach, (iii) gradient-based hierarchical clustering (HC) in hyperbolic space (gHHC) [Monath et al., 2019], (iv) gradient-based Ultrametric Fitting (UltraFit) [Chierchia and Perret, 2019], and (v) HC by hyperbolic Dasgupta’s cost (HHC) [Chami et al., 2020]. We refer to App. B for an extended description of these baselines. For non-TWD baselines, we in addition include the Word Mover’s Distance (WMD) [Kusner et al., 2015] for word-document data and we include the Gene Mover’s Distance (GMD) [Bellazzi et al., 2021] for scRNA-seq data. In the experiments in Sec. 5, the Euclidean distance $\widetilde{d}$ is used as the ground metric for WMD, GMD, and the TWD baselines, which aligns with existing studies [Kusner et al., 2015, Bellazzi et al., 2021, Indyk and Thaper, 2003, Backurs et al., 2020, Le et al., 2019, Chen et al., 2024, Yamada et al., 2022]. For the ablation study that uses cosine distance as the ground metric and directly performs classification tasks in the feature space using Euclidean distance and cosine distance, we refer to App. E.
@@ E.2 Runtime Analysis
In Fig. 4, we present the runtime performance of our TWD in comparison to other TWDs and Wasserstein distances when applied to scRNA-seq and word-document datasets. Although our TWD is not the fastest method, it consistently ranks as the fourth or fifth quickest. It outperforms deep-based methods like HHC-TWD and gHHC-TWD, as well as pre-trained OT baselines such as GMD for single-cell RNA sequencing datasets and WMD for word-document datasets. The fastest methods identified are TSWD-1, TR-TWD, and MST-TWD. The work in [Le et al., 2019] reports that the runtime of TSWD is influenced by the number of tree slices used in sampling. Similarly, slice variants SWCTWD and SWQTWD are also affected by the number of slices. Despite being slightly slower compared to TSWD-1, TR-TWD, and MST-TWD, our TWD demonstrates significant advantages in terms of classification accuracy, as shown in Tab. LABEL:tab:classification. We also report the runtime of the OT distance using the embedding distance between features as a ground distance, denoted as OT-ED. We note that this computational expense is $O(mn^{3}+m^{3}\log m)$. In contrast, our TWD offers a substantially reduced runtime of $O(m^{1.2})$, thereby demonstrating a significant advantage when handling large datasets. Note that the scalability of the proposed TWD is enhanced by utilizing diffusion landmarks [Shen and Wu, 2022], as detailed in App. D.3. This enhancement allows our TWD to achieve computational performance comparable to TSWD-1, TR-TWD, and MST-TWD, particularly in datasets with a large number of features, such as the word-document datasets shown in Fig. 4.
@@ E.3.4 Exploring Other Hyperbolic Embedding and the Impact of Curvature
We note that other TWD baselines could use other two-dimensional Poincaré representations (e.g., Poincaré embedding (PE) [Nickel and Kiela, 2017] or hyperbolic VAE [Nagano et al., 2019]) and decode the hyperbolic embedding into a tree by HHC [Chami et al., 2020]. We denoted these baselines by PE-HHC-TWD and HVAE-HHC-TWD, respectively, and present their results in Tab. E.3.5. We see in the table that HHC [Chami et al., 2020] applied to a PE [Nickel and Kiela, 2017] or hyperbolic VAE [Nagano et al., 2019] is less effective than our TWD by a large margin. This discrepancy is attributed to the sensitivity of the recovery of the hidden feature hierarchy using the competing methods, a finding that is consistent with previous works [Sala et al., 2018, Sonthalia and Gilbert, 2020, Lin et al., 2023].
@@ E.3.5 Robustness Analysis Under Gaussian Noise
PE-HHC-TWD 87.2$\pm$1.2 69.7$\pm$0.9 93.2$\pm$1.1 86.2$\pm$0.9 81.6$\pm$0.9 74.9$\pm$1.0
HVAE-HHC-TWD 88.4$\pm$2.5 64.3$\pm$1.7 90.8$\pm$1.5 87.3$\pm$1.7 80.2$\pm$0.9 75.1$\pm$1.0
E.3.6 Comparison to Non-TWD Baselines
For non-TWD baselines, we also compare our method to Word Mover’s Distance (WMD) [Kusner et al., 2015] for word-document datasets and Gene Mover’s Distance (GMD) [Bellazzi et al., 2021] for scRNA-seq datasets in Sec. 5. Note that WMD [Kusner et al., 2015] computes the Wasserstein distance using the Euclidean distance $\widetilde{d}$ between precomputed Word2Vec embeddings [Mikolov et al., 2013] as the ground cost, while GMD [Bellazzi et al., 2021] is defined similarly, using the Euclidean distance $\widetilde{d}$ between precomputed Gene2Vec embeddings as the ground cost. In the experiments in Sec. 5, the Euclidean distance $\widetilde{d}$ is used as the ground metric for WMD, GMD, and the TWD baselines, which aligns with existing studies [Kusner et al., 2015, Huang et al., 2016, Le et al., 2019, Takezawa et al., 2021, Yamada et al., 2022, Chen et al., 2024, Bellazzi et al., 2021].
We conduct additional experiments evaluating WMD on word-document datasets and GMD on scRNA-seq datasets, both using cosine distance as the ground metric. Tab. E.3.6 presents the document and cell classification accuracy of WMD and GMD using cosine distance as the ground metric. The results show that WMD performs worse with cosine distance as the ground metric compared to Euclidean distance, while GMD performs better with cosine distance than Euclidean distance. However, regardless of the ground metric (cosine or Euclidean), WMD and GMD exhibit less effective performance compared to our method. These findings emphasize the value of inferring latent feature hierarchies, as our approach outperforms the baselines in these tasks.
It is important to note that using the ground metric $\widetilde{d}$ (e.g., Euclidean) in WMD, GMD, or existing TWD methods results in the (approximated) Wasserstein distance being based on the predefined metric $\widetilde{d}$. In contrast, our method employs an initial distance $d$, computed using the cosine similarity in the experiments, to compute an affinity matrix within the diffusion operator, which is central to kernel and manifold learning methods. This process leads to a Wasserstein distance based on a latent hierarchical metric induced by $d$. As noted in App. D.1, the kernel type, scale parameter, and initial distance metric selection in our approach significantly influence its effectiveness. These choices are task-specific and are critical in achieving optimal performance [Lindenbaum et al., 2020a].
Table 5: Document and single-cell classification accuracy of WMD and GMD using cosine distance as the ground metric.
Real-World Word-Document Real-World scRNA-seq
BBCSPORT TWITTER CLASSIC AMAZON Zeisel CBMC
WMD 91.2$\pm$0.4 68.4$\pm$0.2 93.5$\pm$0.3 89.4$\pm$0.2 - -
GMD - - - - 86.9$\pm$0.5 82.7$\pm$0.9
In addition, we conduct additional tests evaluating classification performance directly in the feature space, without using Wasserstein distance or TWD. Tab. E.3.6 presents the document and cell classification accuracy in the feature space using the Euclidean distance and cosine distance. We see that the classification performance relying on generic metrics (Euclidean or cosine) is much less effective than WMD and GMD, which aligns with the findings reported in WMD [Kusner et al., 2015] and GMD [Bellazzi et al., 2021]. This highlights that incorporating the feature relationship into a distance between samples improves the effectiveness of distance metric learning. Furthermore, we see that our TWD method performs better than the distance relying on generic metrics, WMD and GMD. This suggests that inferring explicitly the feature hierarchies is important and leads to the effectiveness of distance metric learning for data with a hierarchical structure.
Table 6: Document and cell classification accuracy in the feature space using the Euclidean distance and cosine distance.
Real-World Word-Document Real-World scRNA-seq
BBCSPORT TWITTER CLASSIC AMAZON Zeisel CBMC
Euclidean 78.2$\pm$1.5 57.3$\pm$0.6 64.0$\pm$0.5 71.3$\pm$0.4 67.2$\pm$1.3 62.5$\pm$0.9
Cosine 91.5$\pm$0.9 68.1$\pm$0.6 90.3$\pm$0.8 86.4$\pm$0.7 73.5$\pm$1.0 72.7$\pm$1.1
@@ F.2 Putting the Features on the Leaves of the Tree
The ground truth latent tree $T$ does not assume that features are positioned at the leaves. In addition, note that the hyperbolic embedding [Lin et al., 2023], which is consistent with other common hyperbolic embedding methods [Chamberlain et al., 2017, Nickel and Kiela, 2017, 2018, Sala et al., 2018], does not restrict all the features to the leaves. Assigning the features to the leaf nodes is only carried out in the tree decoding (Alg. 1), which is in line with other tree decoding algorithms [Sarkar, 2011, Chami et al., 2020, Indyk and Thaper, 2003]. Note that our tree decoding algorithm generates a weighted binary tree. We assert that this does not pose a restriction because any tree can be effectively transformed into a binary tree [Bowditch, 2007, Cormen et al., 2022]. For instance, consider the flexible tree in Fig. 10 (a): [sport[soccer,football,basketball]] (where “sports” is a parent node with three children: “basketball”, “soccer”, and “football”). This tree can be accurately modeled as a (weighted) binary tree as in Fig. 10 (b): [$c$[$b$[$a$[soccer, football], basketball]], sport] (where “sports”, “basketball”, “soccer”, and “football” are all leaves, $a,b,c$ are internal nodes, and the edges encode the same relationships as the flexible tree above).
@@ F.3 Differentiability of Our TWD
    Our TWD is differentiable (details in App. F.3), which is an important property that is lacking in existing TWD baselines.
In addition, we present favorable empirical results in Sec. 5 compared to several TWDs and Wasserstein distances.
@@
F.5 Intuition Behind Single-Linkage Tree Construction in Hyperbolic Embeddings
The single-linkage tree aligns well with hyperbolic embeddings because it inherently emphasizes hierarchical relationships by connecting points based on the measure of similarity determined by their HD-LCAs, where there is the “treelike” nature of a hyperbolic space [Bowditch, 2007]. Specifically, Prop. 4.4 shows that the LCA depth closely approximates the Gromov product, which corresponds to the tree depth in 0-hyperbolic metrics.
In the context of tree-Wasserstein distance (TWD), many tree construction methods aim to approximate a given ground metric $\widetilde{d}$ (often Euclidean) using tree metrics, enabling efficient computation of optimal transport distances. The quality of TWD methods is typically assessed by the mean relative error between the proposed TWD and the Wasserstein distance with the ground metric $\widetilde{d}$. For instance, an ultrametric incurs a $\log m$ distortion with respect to Euclidean distance [Fakcharoenphol et al., 2003], as utilized in UltraTree [Chen et al., 2024]. Hence, the ultrametric Wasserstein distance [Chen et al., 2024] incurs a $\log m$ distortion compared to the original Euclidean Wasserstein distance.
Our approach differs fundamentally from existing TWD methods. Instead of approximating the Wasserstein distance with a given ground metric $\widetilde{d}$ (e.g., Euclidean), we aim to approximate the Wasserstein distance with a latent underlying hierarchical distance via TWD, where we use the tree to represent the latent feature hierarchy.